Fill in the blanks with >, < or = sign.
(a) (-3) + (-6) (-3) – (-6)
(b) (-21) – (-10) (- 31) + (-11)
(c) 45 – (-11) 57 + (-4)
(d) (-25) – (-42) (-42) – (-25)

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) (-3) + (-6) = – [3 + 6] = – 9 and (-3) – (-6) = (-3) + 6 = 3 Here, – 9 < 3 ∴ (- 3) + (- 6) < (- 3) – (- 6) (b) (-21) – (-10) = (-21) + 10 = -11 and (-31) + (-11) = – (31 + 11) = – 42 Here, -42 < -11 or -11 > -42 ∴ (-21) , -(-10) >(-31)+ (-11) (c) 45 – (-11) = 45 + 11 = 56 and 57 + (-4) = 57 -4 = 53 Here, 56 >53 ∴ 45 – (-11) > 57 + ( -4) (d) (-25) – (-42) = -25 + 42 = 17 and (-42) – (-25) = -42 + 25 = -17 Here, 17 > -17 ∴ (-25) – (-42) > (-42) – (-25).